# A Novel Compact LLM Framework for Local, High-Privacy EHR Data Applications

### v1

@@ Abstract
Large Language Models (LLMs) have shown impressive capabilities in natural language processing, yet their use in sensitive domains like healthcare, particularly with Electronic Health Records (EHR), faces significant challenges due to privacy concerns and limited computational resources. This paper presents a compact LLM framework designed for local deployment in settings with strict privacy requirements and limited access to high-performance GPUs. We introduce a novel preprocessing technique that uses information extraction methods, e.g., regular expressions, to filter and emphasize critical information in clinical notes, enhancing the performance of smaller LLMs on EHR data. Our framework is evaluated using zero-shot and few-shot learning paradigms on both private and publicly available (MIMIC-IV) datasets, and we also compare its performance with fine-tuned LLMs on the MIMIC-IV dataset. The results demonstrate that our preprocessing approach significantly boosts the prediction accuracy of smaller LLMs, making them suitable for high-privacy, resource-constrained applications. This study offers valuable insights into optimizing LLM performance for sensitive, data-intensive tasks while addressing computational and privacy limitations.
Keywords LLM  $\cdot$ EHR  $\cdot$ Privacy
@@ 1 Introduction
Large Language Models (LLMs), such as GPT [1] and BERT [2], have gained substantial popularity and usefulness in recent years due to their remarkable ability to understand and generate human-like text. These models leverage vast amounts of data and complex neural network architectures to deliver insights and automation across various fields. The effectiveness of LLMs in processing and producing text has made them invaluable tools in both academic research and industry applications, pushing the boundaries of what machines can achieve in natural language understanding and generation.
Electronic Health Records (EHR) are crucial for modern healthcare as they contain comprehensive patient data [3], including both structured data (e.g., demographics, medications, lab results, genomics) and unstructured data (e.g., clinical notes). The combination enhances personalized clinical decisions, drug discovery, and informative policy-making. Despite the advance in EHR systems, the lack of precise disease phenotyping remains a significant barrier to their full application. While the Internation Classification of Diseases (ICD) codes are recorded in most EHRs, they often fail to capture the details of a patient’s condition [4]. In contrast, clinical notes can more accurately reflect disease progression, but their unstructured nature makes it challenging to extract relevant information. Manual annotation for these notes is labor-intensive, prone to errors [5], and requires highly-skilled healthcare professionals, further complicating the problem. Additionally, the sensitive nature of medical records demands strict adherence to privacy regulations [6], limiting the use of cloud-based LLMs such as ChatGPT.
To tackle the complexities of highly private EHR data, LLMs present a compelling solution to automate annotation processes, reducing the labor and time required to extract disease phenotype. However, the deployment of large-scale LLMs, such as LLaMA-70b, typically requires cloud-based infrastructure, raising privacy concerns due to sensitive health data. In addition, these models demand significant computational power, particularly high-performance GPUs with large amounts of VRAM, which may not be accessible to many labs.
@@
To address these challenges, we propose a novel framework that enhances the performance of smaller LLMs by incorporating a preprocessing step. This step filters and focuses on disease-relevant contexts, reducing the input length and allowing the LLMs to work within their context window limitations. By focusing on the most pertinent sections of the clinical notes, we aim to improve the accuracy of disease phenotyping, while maintaining computational feasibility and data privacy. We present a comparison between our framework and others in Figure 1.
Figure 1: A diagram representing existing methods and our method to deal with EHR dataset.
To evaluate the effectiveness of the proposed method, we apply it to extract disease phenotypes from clinical notes using both zero-shot and few-shot learning paradigms, comparing performance with and without preprocessing. We use metastasis as an example, a critical condition that is not fully captured by ICD codes but is highly important in practice. Additionally, we fine-tune existing LLMs to further assess the impact of our approach. The advantages of our preprocessing method are demonstrated through experimental analysis on two datasets: a private dataset comprising 7,284 head-and-neck cancer (HNC) patients and the publicly available Medical Information Mart for Intensive Care (MIMIC)-IV dataset [9]. Our findings show that preprocessing significantly improves the accuracy of predictions made by smaller LLMs. In summary, the contributions of our work are as follows:
@@
    Introduction of a Preprocessing Framework: We propose a preprocessing framework to filter and highlight key information in long, unstructured clinical notes before feeding into LLMs, enabling smaller LLMs to make more accurate predictions. In this paper, we use regular expressions (regex) as an example.
@@
    Evaluation Across Multiple Learning Paradigms: Our study includes extensive experiments that compare the effectiveness of zero-shot and few-shot learning paradigms, both with and without the application of preprocessing, on EHR data. And we also compare the results with fine-tuned LLMs. This analysis highlights the considerable influence of preprocessing on model performance, enabling researchers with constrained GPU resources to maintain high accuracy while adhering to strict privacy standards.
@@ 2 Related work
Early works of computational phenotyping search keywords from EHRs. Specifically, they first extract disease, medication, or lab-related keywords by simple searching algorithms, and then derive the disease phenotype based on rules designed by domain experts [10]. Rule-based systems, such as Unified Medical Language System [11], were utilized for various phenotypes including diabetes mellitus [12] and adverse drug events [13]. However, the designed rules can fall short in inferring complicated clinical notes. To address this issue, previous studies utilized statistical learning methods to predict disease phenotype after feature engineering the searched keywords and structural data [14, 15, 16]. Although these methods improve the manually designed rules, they usually ignore the context of the searched keywords in the clinical notes and may misinterpret those keywords.
@@
To effectively learn the disease phenotype in the clinical notes, a popular approach is to fine-tune pre-trained language models, such as BERT or clinical BERT [17], using annotated clinical documents. However, fine-tuned models were mainly developed for clinical text mining and named entity extraction [18, 19, 20] and less intended for disease phenotyping, potentially due to the lack of publicly available annotated data. [21] first fine-tuned BERT to extract comprehensive collections of breast cancer-related phenotypes using clinical notes from a Chinese hospital. [22] then adopted a similar approach to develop a cancer domain-specific BERT model (CancerBERT) for breast cancer-related phenotypes extraction from the University of Minnesota Clinical Data Repository (21291 patients). [23] fine-tuned BERT and BART models to automate the generation of discharge summaries for neurology patients, aiming to reduce physician burnout by enhancing the efficiency of clinical documentation. [24] proposed NYUTron, a LLM pre-trained and fine-tuned on inpatient clinical notes of 387,144 patients at New York University to provide an automatic and comprehensive profiling of patients. Notably, all above applications require highly skilled clinicians to annotate notes from a large-scale Private HNC Dataset, which can be labor-intensive and time-consuming.
@@
LLMs can understand contextual language and generate human-like text, excelling across a wide range of NLP tasks [25, 26]. Several studies have directly employed LLMs for various clinical tasks [27]. For example, [28] applied GPT-3.5 to extract social determinants and family history of 1,000 deidentified notes from a University Hospital. [29] employed both GPT-3.5 and GPT-4 to create draft responses for patient messages in electronic inboxes, demonstrating statistically significant decreases in the 4-item physician task load score derivative. [30] employed GPT-4 on private datasets from UW Medical Center Montlake, UW Medical Center Northwest, and Harborview Medical Center to aid clinicians in classifying perioperative risks. [31] introduced a two-agent LLM system named EHR-CoAgent to predict disease from EHR. The system comprises two LLM agents: a predictor agent that makes predictions and articulates reasoning processes, and a critic agent that reviews incorrect predictions and provides feedback for improvements. Remarkably, most of the studies utilized cloud-based LLMs such as OpenAI API instead of locally inferred models, which could result in some privacy issues in more sensitive EHRs.
3 Method
@@ 3.1 Proposed Framework
Our observations indicate that LLMs with fewer parameters often fail to capture essential information when processing extensive text inputs. This issue is compounded in settings where computational resources are constrained, such as laboratories equipped with GPUs with small Video Random Access Memory (VRAM), which limits the deployment of more robust models.
In response to these computational constraints and privacy concerns inherent in processing sensitive data such as EHR, we have developed a structured framework to enhance the analysis of clinical notes.
@@
The initial step involves identifying the most relevant information within the clinical text. In this paper, we demonstrate the use of regular expressions (regex) as a practical example, though it is important to emphasize that other Natural Language Processing (NLP) techniques can also be employed. To illustrate the proposed framework, we focus on the diagnosis of ‘metastasis’ in cancer cohorts. This approach can be readily extended to other diagnoses and patient cohorts with clinical expertise. In diagnosing metastasis, we search for keywords such as ‘metastasis,’ ‘metastatic,’ and other related words and extract the surrounding context for further examination. The full list of “keywords” are listed in the Appendix. This preprocessing step streamlines data handling, reducing the need for labor-intensive manual annotation while enabling more efficient downstream analysis for LLMs. Subsequently, we design customized prompts that are specifically structured to aid the model in accurately identifying medical conditions such as metastasis. These prompts are utilized within both zero-shot and few-shot learning frameworks to refine the model’s predictive capabilities.
The introduction of this preprocessing step markedly enhances the prediction accuracy, a claim that is substantiated by our experiments with the Gemma-7b model, as detailed in this paper. This method not only optimizes the performance of LLMs with fewer parameters but also addresses key operational challenges in handling EHR data within privacy-constrained and resource-limited environments.
3.2 Zero-Shot and Few-Shot learning
@@
To enhance the effectiveness of zero-shot learning, we craft the prompts to guide the large language model’s understanding and analysis of the task at hand. This involves structuring the prompt in a way that directly asks the model to assess whether indications of metastasis are present based on the content of a single clinical note. By doing so, we ensure that the model can make accurate assessments without prior training on specific examples, relying solely on its pre-existing knowledge and the context provided within the prompt.
@@
In our few-shot learning methodology, we adopt a hands-on training approach by carefully selecting a balanced set of examples from the dataset. This collection typically consists of $n$ positive samples, indicating the presence of metastasis, $n$ negative samples, where metastasis is absent, and $n$ neutral samples that lack information relevant to the disease state. We term this selection as $n$-shot learning. These samples are thoughtfully integrated into the learning prompts. This tailored integration is designed to familiarize the large language model with the specific subtleties and requirements of the task at hand. By interacting with this curated set of examples, the model can recalibrate its approach based on the limited yet significant exposure to task-specific data. We list the prompts for both zero-shot and few-shot learning in the Appendix.
3.3 Zero-Shot and Few-Shot learning with preprocessing
The preprocessing of clinical notes starts with the segmentation of the text into individual sentences. This step is essential for the meticulous examination of each sentence on its own, allowing for the precise identification of specific contents within the notes. During this review phase, we rigorously scan each sentence for predefined keywords related to metastasis, including ‘metastasis,’ ‘metastatic,’ and so on, which are listed in the Appendix.
Upon detecting any of these keywords in a sentence, we not only flag that sentence for extraction but also include the sentences immediately before and after it. This approach ensures that we capture a comprehensive context surrounding mentions of metastasis, encompassing any preliminary or subsequent information crucial for an in-depth analysis. Extracting adjacent sentences helps preserve the continuity and relevance of the data, thereby enhancing the accuracy of subsequent interpretations. In this process, we employ regular expressions to identify and extract relevant sentences efficiently. However, more sophisticated NLP techniques could also be utilized to refine this extraction process further.
@@ 3.3 Zero-Shot and Few-Shot learning with preprocessing
Once the preprocessing of the text is completed, we proceed with the analysis using the LLMs. We apply the same prompt that was used in the zero-shot and few-shot learning experiment, which is carefully designed to instruct the model to determine the presence of metastasis based on the preprocessed clinical note.
3.4 LLM fine tuning
We utilize the pre-trained LLMs, Gemma-2b and Gemma-7b, to classify sequences of text based on its extensive training on diverse datasets. Given the absence of definitive ground truth in our dataset, we carefully select 50 samples that explicitly include terms related to metastasis; these are designated as our positive samples. Similarly, we choose 50 samples that suggest the absence of metastatic conditions, which we classify as negative samples.
We then split these 100 curated samples into two sets: 80% of the samples are used for training the model, while the remaining 20% are set aside for validation. This division allows us to train the model on a robust set of data while also holding back a subset to test the model’s ability to generalize to new, unseen data.

### v2

# Enhancing LLMs with Smart Preprocessing for EHR Analysis

@@ Abstract
Large Language Models (LLMs) have demonstrated remarkable proficiency in natural language processing; however, their application in sensitive domains such as healthcare, especially in processing Electronic Health Records (EHRs), is constrained by limited computational resources and privacy concerns. This paper introduces a compact LLM framework optimized for local deployment in environments with stringent privacy requirements and restricted access to high-performance GPUs. Our approach leverages simple yet powerful preprocessing techniques, including regular expressions (regex) and Retrieval-Augmented Generation (RAG), to extract and highlight critical information from clinical notes. By pre-filtering long, unstructured text, we enhance the performance of smaller LLMs on EHR-related tasks. Our framework is evaluated using zero-shot and few-shot learning paradigms on both private and publicly available datasets (MIMIC-IV), with additional comparisons against fine-tuned LLMs on MIMIC-IV. Experimental results demonstrate that our preprocessing strategy significantly supercharges the performance of smaller LLMs, making them well-suited for privacy-sensitive and resource-constrained applications. This study offers valuable insights into optimizing LLM performance for local, secure, and efficient healthcare applications. It provides practical guidance for real-world deployment for LLMs while tackling challenges related to privacy, computational feasibility, and clinical applicability.
Keywords Large Language Models, Privacy-preserving Machine Learning, Data Mining, Local Processing, Electronic Health Records, Regular Expression, Retrieval-Augmented Generation
@@ 1 Introduction
Large Language Models (LLMs), such as GPT [1] and BERT [2], have gained substantial popularity and importance in recent years due to their remarkable ability to understand and generate human-like text. These models leverage vast amounts of data and complex neural network architectures to deliver insights and automation across various fields. The effectiveness of LLMs in processing and producing text has made them invaluable tools in both academic research and industry applications, pushing the boundaries of what machines can achieve in natural language understanding and generation.
Electronic Health Records (EHR) are crucial for modern healthcare as they contain comprehensive patient data [3], including both structured data (e.g., demographics, medications, lab results, genomics) and unstructured data (e.g., clinical notes). The combination enhances personalized clinical decisions, drug discovery, and informative policy-making. Despite the advances in EHR systems, the lack of precise disease phenotyping remains a major challenge for their full utilization. While the International Classification of Diseases (ICD) codes are recorded in most EHRs, they often fail to capture the details of a patient’s condition [4]. In contrast, clinical notes can more accurately reflect disease progression, but their unstructured nature makes it challenging to extract relevant information. Manual annotation for these notes is labor-intensive, prone to errors [5], and requires highly skilled healthcare professionals, further complicating the problem. Additionally, the sensitive nature of medical records demands strict adherence to privacy regulations [6], limiting the use of cloud-based LLMs such as ChatGPT.
To tackle the complexities of the clinical notes in highly private EHR data, LLMs present a compelling solution to automate annotation processes, reducing the labor and time required to extract disease phenotypes. However, the deployment of large-scale LLMs, such as LLaMA-70B, typically requires cloud-based infrastructure, raising privacy concerns due to sensitive health data. In addition, these models demand significant computational power, particularly high-performance GPUs with large amounts of Video Random Access Memory (VRAM), which may not be accessible to many labs.
@@
To address these challenges, we propose a novel framework that enhances the performance of smaller LLMs by incorporating a preprocessing step. This step filters and focuses on disease-relevant contexts, reducing the input length and allowing the LLMs to work within their context window limitations. By focusing on the most pertinent sections of the clinical notes, we aim to improve the accuracy of disease phenotyping, while maintaining computational feasibility and data privacy. We present a comparison between our framework and others in Figure 1. Compared to other methods, our framework incorporates a specifically designed preprocessing step to extract the most relevant information, demonstrated in this paper through regular expressions (regex), which identify predefined relevant keywords, and Retrieval-Augmented Generation (RAG), which retrieves sentences based on semantic similarity when prior knowledge is limited. We should note that these are just two examples to illustrate the power of preprocessing in our framework, and other NLP methods can be easily integrated into the framework.
Figure 1: A diagram representing existing methods and our method to deal with the EHR dataset.
To evaluate the effectiveness of the proposed method, we apply it to extract disease phenotypes from clinical notes using both zero-shot and few-shot learning paradigms, comparing their performance with and without preprocessing. Additionally, we fine-tune existing LLMs to further evaluate the effectiveness of our approach. In this paper, we use “metastasis” as a case study—a critical condition that is not fully captured by ICD codes but holds significant clinical importance. Importantly, our framework can be easily extended to other medical conditions and terminologies, making it broadly applicable across various healthcare domains. The advantages of our preprocessing method are demonstrated through experimental analysis on two datasets: a private dataset comprising 7,284 head-and-neck cancer (HNC) patients and the publicly available Medical Information Mart for Intensive Care (MIMIC)-IV dataset [9]. To further assess stability, we repeat the evaluation on a subset of the MIMIC-IV dataset ten times. Our findings show that preprocessing significantly improves the accuracy of predictions made by smaller LLMs. In summary, the contributions of our work are as follows:
@@
    Introduction of a Preprocessing Framework: We propose a preprocessing framework to filter and highlight key information in long, unstructured clinical notes before feeding into LLMs, enabling smaller LLMs to make more accurate predictions. By leveraging simple yet effective tricks like regex-based filtering and RAG, we demonstrate significant performance improvements in resource-constrained settings. To the best of our knowledge, no previous work has highlighted the effectiveness of simple regex-based approaches in this context.
@@
    Evaluation Across Multiple Learning Paradigms: Our study includes extensive experiments that compare the effectiveness of zero-shot and few-shot learning paradigms, both with and without the application of preprocessing, on EHR data. Different LLMs are also evaluated to ensure robustness. We also compare the results of our framework with fine-tuned LLMs. Experimental results indicate that, in scenarios with limited annotated labels, preprocessing proves to be significantly more efficient and cost-effective than fine-tuning. This analysis highlights the considerable influence of preprocessing on model performance, enabling researchers with constrained GPU resources to maintain high accuracy while adhering to strict privacy standards.
@@ 2 Related Work
Early works of computational phenotyping search keywords from EHRs. Specifically, they first extract disease, medication, or lab-related keywords by simple searching algorithms, and then derive the disease phenotype based on rules designed by domain experts [10]. Rule-based systems, such as The Unified Medical Language System [11], were utilized for various phenotypes including diabetes mellitus [12] and adverse drug events [13]. However, the designed rules can fall short in inferring complicated clinical notes. To address this issue, previous studies utilized statistical learning methods to predict disease phenotype after feature engineering the searched keywords and structural data [14, 15, 16]. Although these methods improve the manually designed rules, they usually ignore the context of the searched keywords in the clinical notes and may misinterpret those keywords.
@@
To effectively learn the disease phenotype in the clinical notes, a popular approach is to fine-tune pre-trained language models, such as BERT or clinical BERT [17], using annotated clinical documents. However, fine-tuned models were mainly developed for clinical text mining and named entity extraction [18, 19, 20] and less intended for disease phenotyping, potentially due to the lack of publicly available annotated data. [21] first fine-tuned BERT to extract comprehensive collections of breast cancer-related phenotypes using clinical notes from a Chinese hospital. [22] then adopted a similar approach to developing a cancer domain-specific BERT model (CancerBERT) for breast cancer-related phenotypes extraction from the University of Minnesota Clinical Data Repository (21291 patients). [23] fine-tuned BERT and BART models to automate the generation of discharge summaries for neurology patients, aiming to reduce physician burnout by enhancing the efficiency of clinical documentation. [24] proposed NYUTron, an LLM pre-trained and fine-tuned on inpatient clinical notes of 387,144 patients at New York University to provide an automatic and comprehensive profiling of patients. Notably, all above applications require highly skilled clinicians to annotate notes from a large-scale Private HNC Dataset, which can be labor-intensive and time-consuming.
@@
LLMs can understand contextual language and generate human-like text, excelling across a wide range of NLP tasks [25, 26]. Several studies have directly employed LLMs for various clinical tasks [27]. For example, [28] applied GPT-3.5 to extract social determinants and family history of 1,000 deidentified notes from a university hospital. [29] employed both GPT-3.5 and GPT-4 to generate draft responses for patient messages in electronic inboxes, demonstrating statistically significant decreases in the 4-item physician task load score derivative. [30] employed GPT-4 on private datasets from UW Medical Center Montlake, UW Medical Center Northwest, and Harborview Medical Center to assist clinicians in classifying perioperative risks. [31] introduced a two-agent LLM system named EHR-CoAgent to predict disease from EHR data. The system comprises two LLM agents: a predictor agent that makes predictions and articulates reasoning processes, and a critic agent that reviews incorrect predictions and provides feedback for improvements. Remarkably, most of the studies utilized cloud-based LLMs such as OpenAI API rather than locally inferred models, which could result in some privacy issues in more sensitive EHRs.
3 Methods
@@ 3.1 Proposed Framework
Our observations indicate that LLMs with fewer parameters often fail to capture essential information when processing extensive text inputs. This issue is compounded in settings where computational resources are constrained, such as laboratories equipped with GPUs with small VRAM, which limits the deployment of more robust models.
@@
In response to these computational constraints and privacy concerns inherent in processing sensitive data such as EHR, we have developed a structured framework to enhance the analysis of clinical notes. The initial step involves identifying the most relevant information within the clinical text. In this paper, we demonstrate the use of regex and RAG as practical examples, though it is important to emphasize that other Natural Language Processing (NLP) techniques can also be employed. To illustrate the proposed framework, we focus on the diagnosis of “metastasis” in cancer cohorts. This approach can be readily extended to other diagnoses and patient cohorts.
First, we emphasize how we use regex to do the preprocessing procedure with clinical expertise. In diagnosing metastasis, we search for keywords such as “metastasis,” “metastatic,” and other related words and extract the surrounding context for further examination. The full list of “keywords” are listed in the Appendix. This preprocessing step streamlines data handling, reducing the need for labor-intensive manual annotation while enabling more efficient downstream analysis for LLMs. Additionally, when prior knowledge is insufficient—such as not knowing the associated terms for “metastasis”—we turn to the RAG approach. RAG combines LLMs with information retrieval systems, retrieving sentences that are semantically similar to input keywords [32] based on cosine similarities. The introduction of these preprocessing steps markedly enhances prediction accuracy, a claim substantiated by our experiments with the Gemma-7B model and other LLMs fine-tuned with biomedical knowledge, as detailed in this paper.
After the preprocessing step, we design customized prompts that are specifically structured to aid the model in accurately identifying medical conditions such as metastasis. These prompts are utilized within both zero-shot and few-shot learning frameworks to refine the model’s predictive capabilities.
Figure 2: Flowchart of our proposed framework and evaluation process for two EHR datasets.
3.2 Evaluations with Zero-Shot and Few-Shot Learning
@@
To evaluate the effectiveness of zero-shot learning, we craft the prompts to guide the large language model’s understanding and analysis of the task at hand. This involves structuring the prompt in a way that directly asks the model to assess whether indications of metastasis are present based on the content of a single clinical note. By doing so, we ensure that the model can make accurate assessments without prior training on specific examples, relying solely on its pre-existing knowledge and the context provided within the prompt.
@@
In our few-shot learning methodology, we adopt a hands-on training approach by carefully selecting a balanced set of examples from the dataset. This collection typically consists of $n$ positive samples, indicating the presence of metastasis, $n$ negative samples, where metastasis is absent, and $n$ neutral samples that lack information relevant to the disease state. We term this selection as “$n$-shot” learning. These samples are thoughtfully integrated into the learning prompts. This integration is designed to familiarize the LLMs with the specific subtleties and requirements of the task at hand. By interacting with this curated set of examples, the model can recalibrate its approach based on the limited yet significant exposure to task-specific data. We list the prompts for both zero-shot and few-shot learning in the Appendix.
We start by processing the entire dataset once using the zero-shot and few-shot learning paradigm to establish overall performance. To further assess stability and variability, we repeat the procedure 10 times using different random seeds, testing only a subset of the MIMIC-IV data due to time constraints. Additionally, we investigate whether varying shot settings improve performance across different LLMs and evaluate output consistency under different conditions. To ensure a comprehensive comparison, we experiment with multiple LLMs, including Gemma-7B, as well as medically fine-tuned models such as Bio-Medical-LLaMA-3-8B and LLaMA-2-7B-Chat-Med.
3.3 Evaluations with Zero-Shot and Few-Shot Learning with Preprocessing
The preprocessing of clinical notes starts with the segmentation of the text into individual sentences. This step is essential for the meticulous examination of each sentence on its own, allowing for the precise identification of specific contents within the notes. After that, we use regex and RAG as two examples to extract the relevant information. Under the regex-based approach, we systematically scan each sentence for predefined keywords related to metastasis, including “metastasis,” “metastatic” and so on, which are listed in the Appendix. Upon detecting any of these keywords in a sentence, we not only flag that sentence for extraction but also include the sentences immediately before and after it. This approach ensures that we capture a comprehensive context surrounding mentions of metastasis, encompassing any preliminary or subsequent information crucial for an in-depth analysis. Extracting adjacent sentences helps preserve the continuity and relevance of the data, thereby enhancing the accuracy of subsequent interpretations. In addition to regex, we utilize the RAG approach when prior knowledge is insufficient or when defining keywords requires domain expertise. Instead of manually determining which sentences are relevant to metastasis, RAG selects and retrieves sentences solely based on similarity scores, measured by Cosine Similarity [33]. This approach provides a viable solution in scenarios where predefined keyword lists are not available.
@@ 3.3 Evaluations with Zero-Shot and Few-Shot Learning with Preprocessing
Once the preprocessing of the text is completed, we investigate whether varying processing settings improve performance on different LLMs and evaluate output consistency under different conditions. We apply the same prompt that was used in the zero-shot and few-shot learning experiment, which is carefully designed to instruct the model to determine the presence of metastasis based on the preprocessed clinical note.
For the regex-based approach, we first process the entire dataset and then conduct 10 additional runs using different random seeds on a subset of the MIMIC-IV dataset to evaluate stability and variability. Additionally, for the RAG-based approach, we repeat the procedure 10 times with different random seeds on the same subset of the MIMIC-IV dataset due to time constraints. Additionally, we assess the impact of extracting varying numbers of relevant sentences in the RAG-based approach to analyze its effect on performance. Similar to the non-preprocessed setting, both approaches are tested across multiple LLMs and different shot settings to ensure consistency and robustness across various models.
3.4 Evaluations with Fine-Tuned LLMs
To compare the effectiveness of preprocessing with the widely used fine-tuning approach, we fine-tune pre-trained LLMs, Gemma-2B and Gemma-7B, for text classification based on their extensive training across diverse datasets. Due to the lack of definitive ground truth in our dataset, we carefully curate 50 samples explicitly containing metastasis-related terms as positive samples. Likewise, we select 50 samples that indicate the absence of metastatic conditions, classifying them as negative samples. We then split these 100 curated samples into two sets: 80% of the samples are used for training the model, while the remaining 20% are set aside for validation. This division allows us to train the model on a robust set of data while also holding back a subset to test the model’s ability to generalize to new, unseen data.